OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Part 1: Determine whether 2 is a zero of the polynomial P(x) = 2x^3 – 5x^2 + 4x – 6 by using the Remainder Theorem. Show your work. (4 points) Part 2: Explain how the Remainder Theorem is useful in finding the zeros of a polynomial function. (4 points) @campbell_st @Loser66 @Mertsj @satellite73 @zepdrix @thomaster

OpenStudy (mertsj):

Replace x with 2 and see if you get 0 for the result.

OpenStudy (magbak):

But can I just get like an explination or something please not just like bam here yo u go now you deal with it.

OpenStudy (mertsj):

magbak...substitute 2 in place of x.

OpenStudy (mertsj):

Simplify. Post the result.

OpenStudy (magbak):

Ok .

OpenStudy (magbak):

But is that using the remainder theorem?

OpenStudy (mertsj):

Ok. You don't have to do that if you don't want to. It is your choice.

OpenStudy (magbak):

trust me I am not trying to be stubborn it is just I want to understand.

OpenStudy (mertsj):

What did you get when you replaced x with 2?

OpenStudy (magbak):

Because I have to use the remainder theorem if I do not then I answered the question wrong even if the answer is right.

OpenStudy (magbak):

Ok give me one sec.

OpenStudy (mertsj):

What did you get when you replaced x with 2?

OpenStudy (magbak):

I got -34

zepdrix (zepdrix):

I think the remainder theorem goes something like this, \(\large x=2 \qquad\to\qquad (x-2)=0\) To use the remainder theorem, we'll want to divide our polynomial by x-2. If we're left with any remainder, then this is not a zero of the function. It's quite a bit more tedious than the process mert is showing you. But if you have to use it, then that's what we should do :)

OpenStudy (magbak):

Is that true @Mertsj and @Loser66

OpenStudy (mertsj):

\[2(2)^3-5(2^2)+4(2)-6=2(8)-5(4)+8-6=16-20+2=-4+2=-2\]

OpenStudy (mertsj):

According to the remainder theorem, since the result is not 0, then 2 is not a zero of the function.

OpenStudy (magbak):

Oh my bad sorry I forgot my PEMDAS exponents first then multiplication.